### v2

# Large Twist Angle Dependent Ultrafast Transient Dynamics of MoSe2/WSe2 van der Waals Heterostructures beyond the Mott Transition

Abstract
Two-dimensional van der Waals heterostructures (HS) exhibit twist-angle () dependent inter-layer charge transfer, driven by moiré potential that tunes the electronic band structure with varying . Apart from the magic angles of 3 ∘ and 57.5 ∘ that show flat valence bands (twisted WSe2 bilayer), the commensurate angles of 21.8∘ and 38.2∘ reveal the Umklapp light coupling of interlayer excitons. We report our results on non-degenerate optical pump-optical probe spectroscopy of MoSe2/WSe2 HS at large twist angles under high photoexcitation densities above the Mott transition threshold, generating interlayer localized charge carriers. We show that the recombination time of electrons and holes is minimum at the commensurate angles. The strength of non-radiative interlayer Auger recombination also shows a minimum at the commensurate angles. The fluence dependence of interlayer carrier recombination time suggests additional relaxation channels near the commensurate angles. This study emphasizes the significance of the large twist angle of HS in developing transition metal dichalcogenides-based optoelectronic devices.
1Department of Physics, Indian Institute of Science, Bangalore 560012, India
2Centre for Ultrafast Laser Applications, Indian Institute of Science, Bangalore 560012, India
32D Materials Research and Innovation Group, Department of Physics, Indian Institute of Technology Madras, Chennai 600036, India
4Centre for Nano and Materials Sciences, Jain (Deemed-to-be University), Jain Global Campus, Kanakpura, Bangalore 562112, Karnataka, India
Keywords: van der Waals Heterostructure, Raman spectroscopy, optical pump-optical probe spectroscopy, ultrafast carrier relaxation dynamics, interlayer excitons, exciton Mott’s transition.
Introduction
Two-dimensional materials such as graphene, transition metal dichalcogenides (TMDs), and h-BN, along with their heterostructures (HS), show many exciting physical properties. For example, in the case of TMDs, properties such as the large binding energy of excitons[1, 2, 3], a strong non-linear optical response[4, 5, 6] and spin-valley coupling[7, 8] are exhibited. For the heterobilayers MoX2/WX2, the energy levels are decided by d orbitals of transition metal. The 5d orbital of W is higher than that of the 4d orbital of Mo, and hence, the conduction band minima and valence band maxima of tungsten are higher than that of molybdenum. This creates a staggered or type-II band alignment[9, 10]. This alignment facilitates ultrafast charge transfer between the layers, generation of interlayer exciton with longer decay times[11, 12], and energy transfer among the layers[13, 3]. The optoelectronic properties have been shown to depend on the twist angle between the layers. The twist between the two layers gives rise to a periodic array of potential minima known as the moiré pattern [14, 15, 16, 17]. At low temperatures and low twist angles, the multiple sharp peaks superimposed over the photoluminescence (PL) spectra denotes the interlayer excitons trapped in moiré quantum wells[18, 19]. Self-organized quantum dots are observed in MoSe2/WSe2 and MoS2/WS2 HS for twist angles less than 0.5∘ [20]. For WS2/MoSe2 HS, twist angle tunes the moiré reciprocal lattice and hence the properties of moiré excitons such as oscillator strength and inter/intralayer mixing [21]. In a twisted WSe2 bilayer, flat top moiré valence bands are observed at magic angles of 3∘ and 57.5∘ [22, 23, 24]. While small twist angles reveal the properties of moiré excitons, large twist angles describe the characteristics of interlayer excitons.
Nayak et al. [10] observed the quenching of photoluminescence for the MoSe2/WSe2 HS for all the twist angles as compared to constituent monolayers. A redshift in PL energy of up to 100 meV has been observed in HS WSe2/MoSe2 depending on the twist angle [25, 26]. It has been reported that the PL intensity is two orders lesser for = 20∘ as compared to = 2∘ for MoSe2/WSe2 HS [27]. The interlayer distance and binding energy of interlayer exciton are twist-dependent and are maximum at 30∘, as revealed by the PL studies [28, 29]. Yu et al. have shown that the interlayer excitons undergo Umklapp recombination near the commensurate twist angles of 21.8∘ and 38.2∘ [27, 30]. The additional channel for recombination of interlayer excitons at these particular angles predicts anomalous relaxation dynamics for the HS. The significance of the commensurate angle 21.8∘ has been emphasized in recent reports on HS MoS2/MoSe2, demonstrating the highest thermoelectric performance (ZT = 2.96 at T = 700 K)[31] and recording the highest indirect bandgap[32].
A recent report on ultrafast electron diffraction of MoSe2/WSe2 HS indicates that interlayer heat transfer occurs on a 20 ps timescale, mediated by nonthermal phonons followed by interlayer charge transfer and scattering[33]. The ultrafast transient absorption spectroscopy of TMDs HS has shown that the interlayer charge transfer happens on the order of 100 fs time scale[34, 35, 36, 37, 38, 39]. A slight variation in geometry or twist angle of the HS can vary the charge transfer time from 100 fs to 1 ps[40, 41, 3]. Zhu et al. have reported the recombination time for interlayer exciton in twisted WSe2/MoS2 HS, varying from 40 ps to 3 ns for twist angles between 0∘ and 30∘, and attributed this variation to non-radiative defect-mediated recombination [42]. In this study, no systematic twist-angle dependence of recombination time was observed. We note that these experiments were performed at a fluence of 1011cm-2, which is below the exciton Mott transition (MT) [42]. The Mott transition occurs above a certain photoexcitation density, where the distance between excitons becomes smaller than their radius, thus screening the Coulomb interactions between electrons and holes resulting in an unbounded electron-hole plasma [43, 44]. The Mott transition of excitons has been observed in MoSe2/WSe2 HS, where interlayer excitons transform into an electron-hole plasma localized within the individual layers at a photoexcitation density of 41012 cm-2, [45]. An abrupt change in the interlayer exciton photoluminescence linewidth has been observed above the MT photoexcitation density [46, 45]. There have been many reports on different heterostructures for carrier relaxation dynamics but primarily for small twist angles [40, 47, 41], with only a few focusing on large twist angles [42]. All these studies are focused on the photoexcitation densities below the MT. This has motivated us to explore the relaxation dynamics above the exciton Mott transition in TMDs HS with large twist angles.
Here, we report the non-degenerate optical pump-optical probe studies of MoSe2/WSe2 HS at different twist angles at ambient conditions. A significant decrease in the recombination time of interlayer electron-hole plasma near the commensurate angles of 21.8∘ and 38.2∘ is observed, which is attributed to additional radiative recombination channels arising from Umklapp recombination. Further, the strength of the interlayer Auger recombination is minimum near the commensurate angles. Our present study can help towards the application of type-II HSs in designing optoelectronic devices, namely photovoltaics, photodetectors, and light-emitting devices where charge transfer and recombination times are important parameters [48].
Experimental details
Non-generate optical pump-optical probe measurements were performed using Ti: Sapphire ultrafast laser system (1 kHz repetition rate, 50 fs pulse width, M/s Spectra-Physics, Spitfire Ace) with the pump and probe energies of 3.1 eV and 1.55 eV, respectively. The pump and probe beams are passed collinearly through an objective lens with the NA = 0.9. The spot size of the beam at the sample point was measured to be 5 m. Experiments were carried out for different twist angles of MoSe2/WSe2 HS at two pump fluences of 2.0 and 2.8 mJ/cm2 at room temperature.
The MoSe2/WSe2 HS were prepared using the chemical vapor deposition method on a c-plane sapphire substrate via a two-step process similar to our earlier work [10]. The individual monolayers of MoSe2 and WSe2 have a triangular geometry as shown in the microscopic image of the sample (Figure 1a). The triangles with light (bright) contrast and bigger (smaller) in size represent WSe2 (MoSe2) monolayers as depicted by the black arrows in Figure 1a. A random alignment of triangular-shaped monolayers was obtained as a result of the CVD process, providing us with various twist angles. It has been verified that the TMDs with triangular geometry terminate with transition metal atoms at the edges using transmission electron microscopy[49] and scanning tunneling microscopy[50]. Therefore, the relative orientation of the layers provides us with the angle between the two layers as there is a direct correlation of crystal orientation with the triangular geometry of each monolayer[40, 49, 51]. The twist angle is evaluated as the angle between two lines collinear to the centroid and the vertex of the triangle[52]. One example of evaluating the twist angle is presented in Figure 1b, where red and blue outlined triangles represent MoSe2 and WSe2 monolayers, respectively. The lines collinear to the vertex and centroid are the angle bisector in the case of equilateral triangles, which gives the twist angle. The schemes for the front and side view of MoSe2/WSe2 HS for different twist angles are shown elsewhere [10].
Raman characterization
The blue and red shifts of the A1g modes in WS2 and MoS2 within their twisted HSs [29] motivated us to perform Raman spectroscopy on HS MoSe2/WSe2 for various twist angles. The group theory for the individual layer provides the irreducible decomposition of Raman active vibration modes A1g(R), E2g(R), and E1g (IR+R) for MX2 monolayer [53]. Figure 2a shows the Raman spectrum of the HS region for a twist angle of 49∘ recorded using 532 nm laser wavelength. We obtain the following vibration modes as shown in Figure 2a: M1(MoSe2, A1g): 238.5 cm-1, M2(WSe2, A1g+E2g): 249 cm-1, M3(WSe2, 2LA): 258.5 cm-1, M4(MoSe2, E2g): 286 cm-1, M5(WSe2, A): 305 cm-1, M6(MoSe2, A): 353 cm-1, M7: 373 cm-1 and M8: 394 cm-1 [54, 55]. It confirms the formation of HS with the presence of all the modes of individual monolayers of MoSe2 and WSe2. From the perspective of the charge transfer among two vertically stacked monolayers, we concentrate on the out-of-plane modes (A1g) of the two monolayers MoSe2 and WSe2 stacked together, namely M1 (MoSe2) and M2 (WSe2).
Figure S1 (supplementary information) shows the normalized intensities of M1, M2, and M3 with different twist angles. The black dashed lines represent the shift of the peak position of modes M1 or M2 at different twist angles. This systematic behavior of M1 and M2 can be seen in Figure 2b, where the dashed red lines (guides to the eyes) show the opposite slopes for M1 and M2 modes for four different ranges of the twist angles. We observed the softening of M1 and hardening of M2 mode for any non-zero twist angle as compared to the aligned geometry (depicted by blue circles for zero twist angle in Figure S1). Mode M2 starts stiffening as the twist angle is varied till 17∘, followed by softening till 32∘ and then stiffening again till 39∘ which is again followed by softening. In a counter behavior, the M1 mode softens as the twist angle is varied till 17∘, followed by stiffening till 32∘ and then softening till 39∘, which is again followed by stiffening. The softening and stiffening of the A1g mode (both M1 and M2) depend upon the carrier doping as well as the interlayer coupling [52, 56, 57]. In transition metal dichalcogenide MoS2, increasing the number of layers stiffens the A1g mode and softens the E2g mode due to increased dielectric screening of Coulomb interactions [58, 59]. It has been reported that the unidirectional transfer of electrons from monolayer WS2 to monolayer MoS2 leads to higher carrier density in MoS2 and lower carrier density in WS2, which can soften or stiffen the respective A1g modes[29]. For the similar staggered bands configuration of HS MoSe2/WSe2, unidirectional electron flow from WSe2 to MoSe2 increases (reduces) the carrier density in MoSe2 (WSe2) and leads to the softening (stiffening) of M1 (M2) modes. To understand the importance of twist angle, we have plotted the difference between the peak position of modes M1 and M2 with twist angles as shown in Figure 2c. A guide to the eye (solid red curve) shows a systematic variation of the shift of M1 with respect to M2 (or vice-versa) with maximum shift near the commensurate angles of 21.8∘ and 38.2∘. For perfectly aligned HS (=0∘), the electron transfer and the interlayer coupling lead to the softening of A1g of both M1 and M2 modes, if we compare these modes in monolayers of MoSe2 and WSe2 (Figure S1). However, with a non-zero twist angle, variation in the charge transfer and the interlayer coupling tunes the softening and the corresponding stiffening of M1 and M2 modes and shows an anomalous maximum shift near the commensurate angles. This can be easily understood as the maximum electron transfer and more dielectric screening of Coulomb interaction between electron and hole for the MoSe2/WSe2 HS at these specific angles, which are in the vicinity of commensurate angles. We also note the disappearance of M3 (2LA mode of WSe2) at the twist angles of 19∘ and 39∘ as depicted by small arrows in Figure S1 (panel iii), which needs further investigation.
Results and discussions
Figure 3a shows time-resolved differential reflectivity (R/R) for monolayer (ML) MoSe2 (blue circles), ML WSe2 (red circles), and HS MoSe2/WSe2 at twist angles () of 20∘ (pink circles) and 35∘ (black circles) with a fluence of 2.0 mJ/cm2. It demonstrates that the relaxation dynamics for the HS differ from those of the individual MLs and vary with different twist angles. For ML MoSe2, a single exponential decay provides a recombination time of 9.60.3 ps, whereas ML WSe2 exhibits a biexponential decay with recombination times of 1.40.15 ps and 20.80.5 ps. A biexponential fit for the R/R(t) (blue circles) for the HS at =20∘ is shown by solid red curve in Figure 3b, using the equation:
where represents the rise part of the time-resolved R/R, with being the rise time, which varies in the range of 0.7-1.0 ps for different twist angles. The fast and slow recombination processes are characterized by amplitudes and recombination times (A1, ) and (A2, ), respectively. A3 is the strength of the slow relaxation process, taken to be constant in the delay time up to 150 ps. Before extracting the fitting parameters for various twist angles, we discuss the mechanism involved in the relaxation dynamics of HS MoSe2/WSe2.
Figure 4 shows the type II band alignment in MoSe2/WSe2 HS depicting higher energy states of valence band maxima (VBM) and conduction band minima (CBM) of WSe2 with respect to that of MoSe2. The optical pump excitation using 400 nm (3.1 eV) laser pulse injects carriers into the conduction band much above the band gap for both constituent monolayers (MoSe2 Eg=1.57 eV, WSe2 Eg=1.67 eV) [60, 61], as shown in the left panel of Figure 4a. The energetic hot carriers undergo rapid electron-electron scattering and relax to the band edge, followed by a charge transfer between the MLs, creating both inter and intralayer excitons [34, 35]. However, at a photoexcitation density of 2 mJ/cm2 ( 6.11014 cm-2), these interlayer excitons undergo Mott transition, where they transform into a hot electron and hot hole plasma in MoSe2 and WSe2 monolayers, respectively [45, 46, 44, 43]. These carriers, localized to individual layers, emerge 1 ps after the photoexcitation, when most of the electrons (holes) get transferred to CBM (VBM) of MoSe2 (WSe2) [45, 35], as shown in Figure 4a. Localized interlayer charge carriers take longer to recombine compared to intralayer carriers, as the electrons and holes are confined to different layers rather than being in the same monolayer. The left panel of Figure 4b represents the radiative recombination of the intralayer electrons and holes in the individual monolayers, associated with (A1, ). For WSe2/MoSe2 HS with non-zero twist angle (42∘), observation of interlayer PL for excitation densities above Mott transition [45, 25] demonstrate the radiative recombination of interlayer electrons and holes, but only at a temperature of 4 K. However, Seyler et al. have reported that interlayer electrons and holes in MoSe2/WSe2 do not align at K valleys in the first Brillouin zone, emphasizing non-radiative recombination as a dominant process [27]. Therefore, we attribute the slow component of the relaxation dynamics (A2, ) to indirect radiative (phonon-mediated) and non-radiative recombination of interlayer electrons and holes, as illustrated in the middle panel of Figure 4b. Motivated by the suggestion of non-radiative Auger recombination of interlayer excitons at high excitation densities [62, 63], the A3 component is attributed to non-radiative interlayer Auger recombination. In this process, electrons and holes from different MLs recombine and transfer the energy to the third particle, such as an electron or a hole, as demonstrated in the right panel of Figure 4b.
Figure 5a shows the twist-angle ()-dependent behavior of the intralayer photoexcited carriers (A1) and its recombination time () at a fluence of 2.0 mJ/cm2. A1 and exhibit no systematic dependence on twist angle as is expected for intralayer carriers recombination in individual monolayers. The recombination dynamics of the interlayer carriers is presented in Figure 5b, where amplitude A2 is twist-angle independent. The recombination time of interlayer localized electrons and holes, (panel ii of Figure 5b) exhibits minima near 20∘ and 38∘. The blue dots represent the data, while the solid blue line serves as a guide to the eye. The component associated with non-radiative Auger recombination of interlayer carriers, A3, also shows minima for 20∘ and 38∘ (Figure 5c). The trend of the fitting parameters at a fluence of 2.0 mJ/cm2 is similar at a fluence of 2.8 mJ/cm2, as shown in supplementary Figure S2.
We offer a qualitative understanding of our results. We observed that is maximum at the twist angle of 30∘ (Figure 5b), i.e. the interlayer localized carriers take the longest time to recombine. This is due to the interlayer distance being maximum at the twist angle of 30∘ in TMDs HSs[28, 29]. Equally interesting is the observation of two minima of observed near 20∘ and 38∘. These twist angles are close to the commensurate angles of = 21.8∘ and 38.2∘ for MoSe2/WSe2 HS, where interlayer localized charge carriers or/and excitons can access an additional channel to recombine[27, 30]. Seyler et al. have demonstrated that for any non-zero twist angle between two layers, the holes in the valence band of WSe2 and electrons in the conduction band of MoSe2 at K points do not align in the first Brillouin zone. However, at the commensurate angle of 21.8∘, the valence band and conduction band align at K points in the second Brillouin zone, as shown in Figure 6, adapted from ref. [27]. At these specific angles, the interlayer charge carriers can recombine radiatively, and the momentum mismatch (due to non-zero twist angle) is compensated by the reciprocal lattice of two layers; the process known as Umklapp recombination [30, 27]. We attribute this Umklapp recombination in the vicinity of commensurate angles as a cause for the two minima observed for (Figure 5b). The two different fluences support the evidence of faster relaxation dynamics in the vicinity of commensurate twist angles. The additional radiative recombination channels for interlayer charge carriers result in lesser interlayer charge carriers involved in non-radiative Auger recombination, leading to a minimum in A3 near these commensurate angles (Figure 5c). To further investigate the minima for near the commensurate angles, the relaxation dynamics is studied as a function of excitation density for different twist angles.
The fluence dependence of the relaxation dynamics of interlayer charge carriers (Figure 5d) shows an increase in interlayer charge carriers with increasing fluence, which saturates beyond 2.0 mJ/cm2 indicating the threshold of charge transfer in HS MoSe2/WSe2 at fluences above 2.0 mJ/cm2. It is important to note that is least fluence dependent at 20∘ and 38∘ in comparison to other twist angles, despite a significant increase in the number of interlayer charge carriers. Intralayer charge carriers do not exhibit this distinct behavior near commensurate angles, as shown by the fluence dependence of amplitude A1 and in supplementary Figure S3. In conclusion, the additional recombination channels for interlayer charge carriers at commensurate angles demonstrate the fluence independence of recombination time () within the reported range up to 2.8 mJ/cm2.
Conclusions
In summary, we have reported the optical pump optical probe spectroscopy of MoSe2/WSe2 HS with different twist angles at high excitation densities above the Mott transition. We observed a significant decrease in interlayer exciton recombination time in the proximity of commensurate angles of 21.8∘ and 38.2∘. We interpret the result as a consequence of the availability of new radiative recombination channels due to Umklapp recombination. The least fluence dependence of interlayer charge carriers recombination time near commensurate angles further supports the presence of additional recombination channels. Our results can be helpful in designing the applications of twisted vdW HS in optoelectronic applications.
Acknowledgments
AKS thanks the Department of Science and Technology, Government of India, for its financial support under the National Science Chair Professorship and Nanomission (Grant No. DST/NM/TUE/QM-5/2019). DVSM also acknowledges the DST for the Nanomission grant. VA acknowledges CSIR for the research fellowship. PKN acknowledges the MHRD STARS research grant [STARS/APR2019/396] and the support from the Institute of Eminence scheme at IIT-Madras through the 2D Materials Research and Innovation Group.

### v3

@@
-Large Twist Angle Dependent Ultrafast Transient Dynamics of MoSe2/WSe2 van der Waals Heterostructures beyond the Mott Transition
+Twist Angle Dependent Ultrafast Transient Dynamics of MoSe2/WSe2 van der Waals Heterostructures beyond the Exciton Mott Transition
 Abstract
@@
 The spot size of the beam at the sample point was measured to be 5 m.
+The spot size is the same as the overlapping region of the heterostructure, and the resulting differential reflectivity () covers the entire overlapping region.
 Experiments were carried out for different twist angles of MoSe2/WSe2 HS at two pump fluences of 2.0 and 2.8 mJ/cm2 at room temperature.
+To investigate carrier relaxation dynamics above the exciton Mott transition, we use a probe beam with a fixed wavelength of 800 nm.
 The MoSe2/WSe2 HS were prepared using the chemical vapor deposition method on a c-plane sapphire substrate via a two-step process similar to our earlier work [10].
+The sapphire substrate were chosen because the strain effects are expected to be minimal (0.08) compared to silica substrate (0.38) [49].
 The individual monolayers of MoSe2 and WSe2 have a triangular geometry as shown in the microscopic image of the sample (Figure 1a).
@@
 From the perspective of the charge transfer among two vertically stacked monolayers, we concentrate on the out-of-plane modes (A1g) of the two monolayers MoSe2 and WSe2 stacked together, namely M1 (MoSe2) and M2 (WSe2).
-Figure S1 (supplementary information) shows the normalized intensities of M1, M2, and M3 with different twist angles.
+Figure S1 (supplementary information) shows the Raman spectra of M1, M2, and M3 modes with normalized intensities for different twist angles.
 The black dashed lines represent the shift of the peak position of modes M1 or M2 at different twist angles.
-This systematic behavior of M1 and M2 can be seen in Figure 2b, where the dashed red lines (guides to the eyes) show the opposite slopes for M1 and M2 modes for four different ranges of the twist angles.
-We observed the softening of M1 and hardening of M2 mode for any non-zero twist angle as compared to the aligned geometry (depicted by blue circles for zero twist angle in Figure S1).
+We observed the softening of M1 and hardening of M2 mode for any non-zero twist angle as compared to the aligned geometry.
 Mode M2 starts stiffening as the twist angle is varied till 17∘, followed by softening till 32∘ and then stiffening again till 39∘ which is again followed by softening.
@@
 For the similar staggered bands configuration of HS MoSe2/WSe2, unidirectional electron flow from WSe2 to MoSe2 increases (reduces) the carrier density in MoSe2 (WSe2) and leads to the softening (stiffening) of M1 (M2) modes.
-To understand the importance of twist angle, we have plotted the difference between the peak position of modes M1 and M2 with twist angles as shown in Figure 2c.
-A guide to the eye (solid red curve) shows a systematic variation of the shift of M1 with respect to M2 (or vice-versa) with maximum shift near the commensurate angles of 21.8∘ and 38.2∘.
+To understand the importance of twist angle, we have plotted the difference between the peak position of modes M1 and M2 with twist angles as shown in Figure 2b.
+Here, the solid blue line shows a systematic variation of the shift of M1 with respect to M2 (or vice-versa) with maximum shift near the commensurate angles of 21.8∘ and 38.2∘.
 For perfectly aligned HS (=0∘), the electron transfer and the interlayer coupling lead to the softening of A1g of both M1 and M2 modes, if we compare these modes in monolayers of MoSe2 and WSe2 (Figure S1).
@@
 Before extracting the fitting parameters for various twist angles, we discuss the mechanism involved in the relaxation dynamics of HS MoSe2/WSe2.
-Figure 4 shows the type II band alignment in MoSe2/WSe2 HS depicting higher energy states of valence band maxima (VBM) and conduction band minima (CBM) of WSe2 with respect to that of MoSe2.
-The optical pump excitation using 400 nm (3.1 eV) laser pulse injects carriers into the conduction band much above the band gap for both constituent monolayers (MoSe2 Eg=1.57 eV, WSe2 Eg=1.67 eV) [60, 61], as shown in the left panel of Figure 4a.
+Figure 4a shows the type II band alignment in MoSe2/WSe2 HS depicting higher energy states of valence band maxima (VBM) and conduction band minima (CBM) of WSe2 with respect to that of MoSe2.
+The optical pump excitation using 400 nm (3.1 eV) laser pulse injects carriers into the conduction band much above the band gap for both constituent monolayers (MoSe2 Eg=1.57 eV, WSe2 Eg=1.67 eV) [61, 62], as shown in the top left panel of Figure 4a.
 The energetic hot carriers undergo rapid electron-electron scattering and relax to the band edge, followed by a charge transfer between the MLs, creating both inter and intralayer excitons [34, 35].
 However, at a photoexcitation density of 2 mJ/cm2 ( 6.11014 cm-2), these interlayer excitons undergo Mott transition, where they transform into a hot electron and hot hole plasma in MoSe2 and WSe2 monolayers, respectively [45, 46, 44, 43].
-These carriers, localized to individual layers, emerge 1 ps after the photoexcitation, when most of the electrons (holes) get transferred to CBM (VBM) of MoSe2 (WSe2) [45, 35], as shown in Figure 4a.
+These carriers, localized to individual layers, emerge 1 ps after the photoexcitation, when most of the electrons (holes) get transferred to CBM (VBM) of MoSe2 (WSe2) [45, 35], as shown in the top right panel of Figure 4a.
 Localized interlayer charge carriers take longer to recombine compared to intralayer carriers, as the electrons and holes are confined to different layers rather than being in the same monolayer.
-The left panel of Figure 4b represents the radiative recombination of the intralayer electrons and holes in the individual monolayers, associated with (A1, ).
+The bottom left panel of Figure 4a represents the radiative recombination of the intralayer electrons and holes in the individual monolayers, associated with (A1, ).
 For WSe2/MoSe2 HS with non-zero twist angle (42∘), observation of interlayer PL for excitation densities above Mott transition [45, 25] demonstrate the radiative recombination of interlayer electrons and holes, but only at a temperature of 4 K.
@@
 have reported that interlayer electrons and holes in MoSe2/WSe2 do not align at K valleys in the first Brillouin zone, emphasizing non-radiative recombination as a dominant process [27].
-Therefore, we attribute the slow component of the relaxation dynamics (A2, ) to indirect radiative (phonon-mediated) and non-radiative recombination of interlayer electrons and holes, as illustrated in the middle panel of Figure 4b.
+Therefore, we attribute the slow component of the relaxation dynamics (A2, ) to indirect radiative (phonon-mediated) and non-radiative recombination of interlayer electrons and holes, as illustrated in the lower middle panel of Figure 4a.
 Motivated by the suggestion of non-radiative Auger recombination of interlayer excitons at high excitation densities [63, 64], the A3 component is attributed to non-radiative interlayer Auger recombination.
-In this process, electrons and holes from different MLs recombine and transfer the energy to the third particle, such as an electron or a hole, as demonstrated in the right panel of Figure 4b.
+In this process, electrons and holes from different MLs recombine and transfer the energy to the third particle, such as an electron or a hole, as demonstrated in the lower right panel of Figure 4a.
 Figure 5a shows the twist-angle ()-dependent behavior of the intralayer photoexcited carriers (A1) and its recombination time () at a fluence of 2.0 mJ/cm2.
@@
 The recombination time of interlayer localized electrons and holes, (panel ii of Figure 5b) exhibits minima near 20∘ and 38∘.
-The blue dots represent the data, while the solid blue line serves as a guide to the eye.
-The component associated with non-radiative Auger recombination of interlayer carriers, A3, also shows minima for 20∘ and 38∘ (Figure 5c).
+The blue dots represent the data, while the solid blue line is a cubic spline.
+The component associated with non-radiative Auger recombination of interlayer carriers, A3, also shows minima for 20∘ and 38∘ (represented by the solid blue line in Figure 5c).
 The trend of the fitting parameters at a fluence of 2.0 mJ/cm2 is similar at a fluence of 2.8 mJ/cm2, as shown in supplementary Figure S2.
@@
 have demonstrated that for any non-zero twist angle between two layers, the holes in the valence band of WSe2 and electrons in the conduction band of MoSe2 at K points do not align in the first Brillouin zone.
-However, at the commensurate angle of 21.8∘, the valence band and conduction band align at K points in the second Brillouin zone, as shown in Figure 6, adapted from ref.
+However, at the commensurate angle of 21.8∘, the valence band and conduction band align at K points in the second Brillouin zone, as shown in Figure 4b, adapted from ref.
 [27].
@@
 The two different fluences support the evidence of faster relaxation dynamics in the vicinity of commensurate twist angles.
-The additional radiative recombination channels for interlayer charge carriers result in lesser interlayer charge carriers involved in non-radiative Auger recombination, leading to a minimum in A3 near these commensurate angles (Figure 5c).
+The additional radiative recombination channels for interlayer charge carriers result in lesser interlayer charge carriers involved in non-radiative Auger recombination, leading to a trend that shows minima in A3 near these commensurate angles (Figure 5c).
 To further investigate the minima for near the commensurate angles, the relaxation dynamics is studied as a function of excitation density for different twist angles.
@@
 PKN acknowledges the MHRD STARS research grant [STARS/APR2019/396] and the support from the Institute of Eminence scheme at IIT-Madras through the 2D Materials Research and Innovation Group.
+[figure]style=plain,subcapbesideposition=top